OpenStudy (anonymous):

If f(x) =x+3/3, then what is f-1(x) ?

OpenStudy (anonymous):

f^-1

OpenStudy (e.mccormick):

OK. Know the process for finding the inverse?

OpenStudy (anonymous):

i dont think so

OpenStudy (e.mccormick):

Fist, f(x) = y. Second, swap the x and y. Third, solve for x. The new equation will be the inverse.

OpenStudy (e.mccormick):

\[f(x)=\frac{ x+3 }{3 } \rightarrow y=\frac{ x+3 }{3 } \rightarrow x=\frac{ y+3 }{3 }\] that is the start of it.

OpenStudy (e.mccormick):

Now, solve for y. I said X earlier, and I meant y. Sorry.